Find the total translational kinetic energy of
3 L of oxygen gas held at a temperature of
−3C◦ and a pressure of 2 atm.
Answer in units of J.
1
Expert's answer
2020-03-30T08:02:03-0400

Solution:

For every degree of freedom of the molecule gas has the same average energy


"\varepsilon_i=\frac {1}{2} kT"

where k is the Boltzmann constant, T is the thermodynamic temperature of the gas. The translational motion of a diatomic oxygen molecule corresponds to three degrees of freedom, and rotational - two. Then the average kinetic energy of the molecule


"\varepsilon = \frac {5}{2} kT"

The kinetic energy of motion of all gas molecules


"E=N \varepsilon"

The number of all gas molecules N, concentration of molecules n


"p=nkT"


"n=\frac {N}{V}"

"N= \frac {pV}{kT}"


"E= N \times \frac {5}{2} kT"

"E= \frac {pV}{kT}\times \frac {5}{2} kT"


"E= \frac {5}{2} pV"

"p=2\times 101325.011 \frac {N}{m^2}= 2.03 \times 10^5 \frac {N}{m^2}"

"V=3\times 10^{-3} m^3"


"E=1.5\times 10^3 J"

Answer: "1.5\times 10^3 J"